# High-Dimensional Gaussian Process Regression with Soft Kernel Interpolation

### v2

@@ 3 Related Work
Recent work on improving the scaling properties of SKI has been focused on efficient incorporation of simplicial interpolation, which when implemented carefully in the case of Simplex-GP (Kapoor et al., 2021) and Sparse grid GP (Yadav et al., 2023) can reduce the cost of a single MVM with $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ down to $\mathcal{O}\left(nd^{2}\right)$. These methods differ in the grid construction used for interpolation with Simplex-SKI opting for a permutohedral lattice (Adams et al., 2010), while sparse grid-GP uses sparse grids and a bespoke recursive MVM algorithm, leading to $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ MVM costs of $\mathcal{O}(d^{2}(n+m))$ and $\mathcal{O}\left(m(\log m)^{d}\right)$ respectively. While both of these methods greatly improve the scaling of SKI, they still feature a problematic dependency on $d$ limiting their extension to arbitrarily high dimensional datasets.
4 Method
@@
Figure 1: (Structured & Soft Kernel Interpolation): Comparison of interpolation procedure for local cubic interpolation during SKI and softmax interpolation during SoftKI. Here white points $\circ$ indicate interpolation points $z_{i}\in\mathbf{z}$ and black points $\bullet$ are data points $x_{i}\in\mathbf{x}$.
4.1 Soft Kernel Interpolation
@@
Define softmax interpolation weights as the following matrix
$\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}=\left[\frac{\exp\left(-\left\|x_{i}-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x_{i}-z_{k}\right\|\right)}\right]_{ij}$ (10)
to create the SoftKI kernel
$\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}=\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{\Sigma}_{\mathbf{z}\mathbf{x}}\,.$ (11)
@@
$\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}=\left[\frac{\exp\left(-\left\|x_{i}/T-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x_{i}/T-z_{k}\right\|\right)}\right]_{ij}\,.$ (12)
@@
This additional hyperparameter is needed because unlike the SGPR kernel $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}=\mathbf{K}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}$ where the lengthscale influences each term, the lengthscale only affects $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ in the SoftKI kernel $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}=\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{\Sigma}_{\mathbf{z}\mathbf{x}}$. The temperature acts as lengthscale-like hyperparameter on the interpolation scheme that controls the distance between the datapoints $\mathbf{x}$ and the interpolation points $\mathbf{z}$. When $T=1$, we obtain the original scheme give in Equation 10. Analogous to how automatic relevance detection (ARD) (MacKay et al., 1994) can be used to set lengthscales for different dimensions, we can also use a different temperature per dimension. When learning both temperatures and lengthscales, we cap the lengthscale range for more stable hyperparameter optimization since they have a push-and-pull effect.
@@
5$\triangleright$ Model Training
6:$\mathbf{z}\leftarrow\texttt{kmeans}(\mathbf{x},m)$
@@
12:         $\mathbf{\theta}\leftarrow\mathbf{\theta}+\eta\,\nabla_{\mathbf{\theta}}\log\widehat{p}\Big{(}\mathbf{y}_{b}\,\big{|}\,\mathbf{x}_{b};\,\mathbf{K}_{\mathbf{x}_{b}\mathbf{x}_{b}}^{\text{SoftKI}}+\mathbf{\Lambda}\Big{)}$triangleright Stabilized MLL (Section 4.2)
@@
15$\triangleright$ Stabilized Inference (Section 4.3
16:$\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\leftarrow\texttt{softmax\_interpolation}(\mathbf{x},\mathbf{z})$
17:$\mathbf{U}_{\mathbf{z}\mathbf{z}}^{\top}\mathbf{U}_{\mathbf{z}\mathbf{z}}\leftarrow\texttt{cholesky}(\mathbf{K}_{\mathbf{z}\mathbf{z}})$
18:$\mathbf{Q},\mathbf{R}\leftarrow\texttt{QR}\!\left(\begin{pmatrix}\mathbf{\Lambda}^{-1/2}\,\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\\[4.30554pt] \mathbf{U}_{\mathbf{z}\mathbf{z}}\end{pmatrix}\right)$
19:$\alpha\leftarrow\mathbf{R}^{-1}\,\mathbf{Q}^{\top}\!\begin{pmatrix}\mathbf{\Lambda}^{-1/2}\,\mathbf{y}\\[4.30554pt] 0\end{pmatrix}$
@@
Figure 2: (Inducing vs. Interpolation Points): Comparison of inducing points learned by SGPR, static rectilinear lattice points in SKI, and adaptive interpolation points learned by SoftKI on the Ricker wavelet  (Ricker, 1953) after 100 epochs of hyperparameter optimization. Contour plots of the posterior mean are paired with each method’s absolute-error. Each method achieves comparable accuracy: SGPR’s inducing points follow the sample distribution, and SoftKI’s interpolation points adapt to the true function’s local geometry. Additional experiment details can be found in Appendix C.1
4.2 Learning Interpolation Points
with derivative
$\frac{\partial\log p(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})}{\partial\mathbf{\theta}}$ $=-\frac{1}{2}\Big{[}\mathbf{y}^{\top}\mathbf{D}_{\theta}^{-1}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{D}_{\theta}^{-1}\mathbf{y}+\operatorname{tr}\Big{(}\mathbf{D}_{\theta}^{-1}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\Big{)}\Big{]}.$ (14)
Each evaluation of $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ has time complexity $\mathcal{O}(m^{2}n)$ and space complexity $\mathcal{O}(mn)$ where $m$ is the number of interpolation points. This is tractable to compute only when $m\ll n$ because $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}$ is low-rank.
Hutchinson’s pseudoloss. Occasionally, we observe that the MLL is numerically unstable in single-precision floating point arithmetic (see Appendix B.1 for further details). To help stabilize the MLL in these situations, we use an approximate MLL based on work in approximate GP theory (Gardner et al., 2018; Maddox et al., 2021; Wenger et al., 2023) that identify efficiently computable estimates of the log determinant (Equation 13) and trace term (Equation 14) in the GP MLL. In more detail, these methods combine stochastic trace estimation via the Hutchinson’s trace estimator (Girard, 1989; Hutchinson, 1989) with blocked conjugate gradients so that the overall cost remains quadratic in the size of the kernel. Maddox et al. (2022) term this approximate MLL the Hutchinson’s pseudoloss and show that it remains stable under low-precision conditions.
Definition 4.2 (Hutchinson Pseudoloss).
Let $\mathbf{u}_{0},\mathbf{u}_{1},\dots,\mathbf{u}_{l}$ be the solutions obtained by using block conjugate gradients for the system $\mathbf{D}_{\mathbf{\theta}}\bigl{(}\mathbf{u}_{0}\;\mathbf{u}_{1}\dots\mathbf{u}_{l}\bigr{)}=\bigl{(}\mathbf{y}\;\mathbf{z}_{1}\dots\mathbf{z}_{l}\bigr{)},$ where each $\mathbf{z}_{j}$ is a Gaussian random vector normalized to unit length. The Hutchinson pseudoloss approximation of the GP marginal log likelihood in Equation 13 is given as
$\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})=-\frac{1}{2}\left[\mathbf{u}_{0}^{\top}\mathbf{D}_{\mathbf{\theta}}\mathbf{u}_{0}+\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}(\mathbf{D}_{\mathbf{\theta}}\mathbf{z}_{j})\right]$ (15)
with derivative
$\frac{\partial\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})}{\partial\mathbf{\theta}}=-\frac{1}{2}\left[\mathbf{u}_{0}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{u}_{0}+\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{z}_{j}\right]$ (16)
By computing Equation 15 as described above as opposed to explicitly computing a stochastic Lanczos quadrature approximation (Ubaru et al., 2017), the gradient can also be efficiently computed using back propagation.
$\log\hat{p}(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})=\begin{cases}\log p(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})&\mbox{when stable}\\ \log\tilde{p}(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})&\mbox{otherwise}\end{cases}$ (17)
for optimizing SoftKI’s hyperparameters. In practice, this means that we default to using SoftKI’s MLL and fallback to Hutchinson’s pseudoloss when numeric instability is encountered. This enables SoftKI to accurately recover approximations of the MLL, even when the current positioning of interpolation points would otherwise make the direct computation of the MLL unstable, while relying on the exact MLL as much as possible.
Stochastic optimization. Instead of computing $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ on the entire dataset $\mathbf{x}$, we can compute $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x}_{b};\mathbf{\theta})$ on a minibatch of data $\mathbf{x}_{b}$ of size $b$ to perform stochastic optimization (see line 7 of Algorithm 1). The stochastic estimate provides an unbiased estimator of the gradient. In this way, SoftKI can leverage powerful stochastic optimization techniques used to train neural networks such as Adam (Kingma & Ba, 2014) and hardware acceleration such as graphics processing units (GPUs). The time complexity of evaluating $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x}_{b};\mathbf{\theta})$ is cheap, costing $\mathcal{O}(b^{2}m)$ operations with space complexity $\mathcal{O}(bm)$. The lower space complexity makes it easier to use GPUs since they have more memory constraints compared to CPUs. We emphasize that we are performing stochastic gradient descent and not stochastic variational inference as in SVGP (Hensman et al., 2013). In particular, we do not define a distribution on the interpolation points $\mathbf{z}$ nor make a variational approximation.
@@ 4.3 Posterior Inference
$p(f(*)\,|\,\mathbf{y})=\mathcal{N}(\mathbf{\widehat{K}_{*\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{y},\mathbf{K}_{**}^{\text{SoftKI}}-\mathbf{K}_{*\mathbf{x}}^{\text{SoftKI}}(\mathbf{\Lambda}^{-1}-\mathbf{\Lambda}^{-1}\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1})\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}})$ (18)
@@
for weights $\alpha$. Note that this is the posterior mean of a SGPR with $\mathbf{C}$ replaced with $\widehat{\mathbf{C}}$ and $\mathbf{K_{\mathbf{x}\mathbf{z}}}$ replaced with $\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}$. Moreover, note that $\mathbf{\widehat{C}}$ is $\mathbf{C}$ with $\mathbf{K_{\mathbf{x}\mathbf{z}}}$ replaced with $\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}$. Since $\mathbf{\widehat{C}}$ is a $m\times m$ matrix, the solution of the system of linear equations has time complexity $\mathcal{O}(m^{3})$. The formation of $\mathbf{\widehat{C}}$ requires the multiplication of a $m\times n$ matrix with a $n\times m$ matrix which has time complexity $\mathcal{O}(m^{2}n)$. Thus the complexity of SoftKI posterior mean inference is $\mathcal{O}(m^{2}n)$ since it is dominated by the formation of $\mathbf{\widehat{C}}$. We refer the reader to the supplementary material for discussion of the posterior covariance (Appendix A).
Solving with QR. Unfortunately, solving Equation 19 for $\alpha$ can be numerically unstable. Foster et al. (2009) introduce a stable QR solver approach for a Subset of Regressors (SoR) GP (Smola & Bartlett, 2000) which we adapt to a SoftKI. Define the block matrix
$\mathbf{A}=\begin{pmatrix}\mathbf{\Lambda}^{-1/2}\mathbf{\widehat{K}}_{\mathbf{x}\mathbf{z}}\\ \mathbf{U}_{\mathbf{z}\mathbf{z}}\end{pmatrix}$ (20)
@@
$\mathbf{A}^{\top}\mathbf{A}$ $=\mathbf{\widehat{K}}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda}^{-1}\mathbf{\widehat{K}}_{\mathbf{x}\mathbf{z}}+\mathbf{K}_{\mathbf{z}\mathbf{z}}=\mathbf{\widehat{C}}\,.$ (21)
@@
We thus solve Equation 22 for $\alpha$ via a triangular solve in $\mathcal{O}(m^{2})$ time. For additional experiments demonstrating the inability of other linear solvers to offer comparable accuracy to the QR-stabilized solve detailed in this section, see Appendix B.2.

### v3

@@ 2.3 Structured Kernel Interpolation
Recent work on improving the scaling properties of SKI has been focused on efficient incorporation of simplicial interpolation, which when implemented carefully in the case of Simplex-GP (Kapoor et al., 2021) and Sparse-Grid GP (Yadav et al., 2022) can reduce the cost of a single MVM with $\mathbf{W}_{\mathbf{x}\mathbf{z}}$ down to $\mathcal{O}\left(nd^{2}\right)$. These methods differ in the grid construction used for interpolation with Simplex-SKI opting for a permutohedral lattice (Adams et al., 2010), while Sparse-Grid GP uses sparse grids (Bungartz & Griebel, 2004) and a custom recursive MVM algorithm. These modifications vastly improve KISS-GP leading to $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ MVM costs of $\mathcal{O}(d^{2}(n+m))$ for Simplex-GP and $\mathcal{O}\left(nd^{2+\ell}+\ell^{d}2^{\ell}\right)$ for Sparse-Grid GP where $\ell$ is a small constant. While both of these methods greatly improve the scaling of SKI, they still feature a problematic dependency on $d$ limiting their extension to arbitrarily high dimensional datasets. The complexity of SKI and related variants are tabulated in LABEL:table:SKI.
3 Soft Kernel Interpolation
@@
3.1 Soft Interpolation Strategies
@@
For a finite collection of inputs $\{x_{i}\}_{i=1}^{n}$, $x_{i}\in\mathbb{R}^{d}$, and a set of interpolation points $\{z_{j}\}_{j=1}^{m}$, $z_{j}\in\mathbb{R}^{d}$, define the softmax interpolation weight matrix as
$\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}=\left[\frac{\exp\left(-\left\|x_{i}-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x_{i}-z_{k}\right\|\right)}\right]_{ij}.$ (4)
With this definition in hand we can now define the SoftKI kernel approximation
$\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}=\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}\mathbf{\Sigma}_{\mathbf{z}\mathbf{x}}\approx\mathbf{K}_{\mathbf{x}\mathbf{x}}.$
@@
$\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}=\left[\frac{\exp\left(-\left\|x_{i}/T-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x_{i}/T-z_{k}\right\|\right)}\right]_{ij}\,.$
@@
This additional hyperparameter is needed because unlike the SGPR kernel $\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SGPR}}=\mathbf{K}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}^{-1}\mathbf{K}_{\mathbf{z}\mathbf{x}}$ where the lengthscale influences each term, the lengthscale only affects $\mathbf{K}_{\mathbf{z}\mathbf{z}}$ in the SoftKI kernel. The temperature acts as lengthscale-like hyperparameter on the interpolation scheme that controls the distance between the datapoints $\mathbf{x}$ and the interpolation points $\mathbf{z}$. When $T=1$, we obtain the original scheme give in Equation 4. Analogous to how automatic relevance detection (ARD) (MacKay et al., 1994) can be used to set lengthscales for different dimensions, we can also use a different temperature per dimension. When learning both temperatures and lengthscales, we cap the lengthscale range for more stable hyperparameter optimization since they have a push-and-pull effect.
Figure 1: Structured & Soft Kernel Interpolation: Comparison of interpolation procedure for local cubic interpolation during KISS-GP (left) and softmax interpolation (right) during SoftKI. Here white points $\circ$ indicate interpolation points $z_{i}\in\mathbf{z}$ and black points $\bullet$ are data points $x_{i}\in\mathbf{x}$. In these diagrams magenta line segments indicate which interpolation points are being used for a given data point during the interpolation procedure of each method. In KISS-GP, a small local subset of grid points is used, whereas in SoftKI all interpolation points are included, weighted by a softmax distribution centered at a given data point.
3.2 Learning Interpolation Points
and derivative
$\frac{\partial\log p(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})}{\partial\mathbf{\theta}}$ $=-\frac{1}{2}\Big{[}\mathbf{y}^{\top}\mathbf{D}_{\theta}^{-1}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{D}_{\theta}^{-1}\mathbf{y}-\operatorname{tr}\Big{(}\mathbf{D}_{\theta}^{-1}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\Big{)}\Big{]}.$ (6)
Each evaluation of $\log p(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ has time complexity $\mathcal{O}(nm^{2})$ and space complexity $\mathcal{O}(nm)$ where $m$ is the number of interpolation points.
Hutchinson’s pseudoloss. Occasionally, we observe that the MLL is numerically unstable in single-precision floating point arithmetic (see Section 4.3 for further details). To help stabilize the MLL in these situations, we use an approximate MLL based on work in approximate GP theory (Gardner et al., 2018b; Maddox et al., 2021; Wenger et al., 2023) that identify efficiently computable estimates of the log determinant (Equation 5) and trace term (Equation 6) in the GP MLL. In more detail, these methods combine stochastic trace estimation via the Hutchinson’s trace estimator (Girard, 1989; Hutchinson, 1989) with blocked CG so that the overall cost remains quadratic in the size of the kernel. Maddox et al. (2022) term this approximate MLL the Hutchinson’s pseudoloss and show that it remains stable under low-precision conditions.
Figure 2: Inducing vs. Interpolation Points: Comparison of inducing points learned by SGPR, static rectilinear lattice points in SKI, and adaptive interpolation points learned by SoftKI on the Ricker wavelet  (Ricker, 1953) after 100 epochs of hyperparameter optimization. Contour plots of the posterior mean are paired with each method’s absolute-error. Each method achieves comparable accuracy: SGPR’s inducing points follow the sample distribution, and SoftKI’s interpolation points adapt to the true function’s local geometry. Additional experiment details can be found in Appendix C.1
Definition 3.2 (Hutchinson Pseudoloss).
Let $\mathbf{u}_{0},\mathbf{u}_{1},\dots,\mathbf{u}_{l}$ be the solutions obtained by using block conjugate gradients for the system $\mathbf{D}_{\mathbf{\theta}}\bigl{(}\mathbf{u}_{0}\;\mathbf{u}_{1}\dots\mathbf{u}_{l}\bigr{)}=\bigl{(}\mathbf{y}\;\mathbf{w}_{1}\dots\mathbf{w}_{l}\bigr{)},$ where each $\mathbf{w}_{j}\in\mathbb{R}^{n}$ is a Gaussian random vector normalized to unit length. The Hutchinson pseudoloss approximation of the GP marginal log likelihood in Equation 5 is given as
$\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})=-\frac{1}{2}\Big{[}\mathbf{u}_{0}^{\top}\mathbf{D}_{\mathbf{\theta}}\mathbf{u}_{0}+\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}(\mathbf{D}_{\mathbf{\theta}}\mathbf{w}_{j})\Big{]}$ (7)
with derivative
$\frac{\partial\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})}{\partial\mathbf{\theta}}=-\frac{1}{2}\Big{[}\mathbf{u}_{0}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{u}_{0}-\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{w}_{j}\Big{]}$ (8)
By computing Equation 7 as described above as opposed to explicitly computing a stochastic Lanczos quadrature approximation (Ubaru et al., 2017), the gradient can also be efficiently computed using back propagation. Note that the gradient of Hutchinson’s pseudoloss approximates the gradient of the GP MLL, where the trace term is estimated using a stochastic trace estimator. The accuracy of this approximation depends on the number of probe vectors $\ell$. Consequently, gradient-based optimization of the GP MLL can be approximated by optimizing Hutchinson’s pseudoloss instead.
$\log\hat{p}(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})=\begin{cases}\log p(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})&\mbox{when stable}\\ \log\tilde{p}(\mathbf{y}\mid\mathbf{x};\mathbf{\theta})&\mbox{otherwise}\end{cases}$
for optimizing SoftKI’s hyperparameters. In practice, this means that we default to using SoftKI’s MLL and fallback to Hutchinson’s pseudoloss when numeric instability is encountered. This enables SoftKI to accurately recover approximations of the MLL’s gradient during gradient-based optimization, even when the current positioning of interpolation points would otherwise make the direct computation of the MLL’s gradient unstable, while relying on the exact MLL as much as possible.
Stochastic optimization. Instead of computing $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x};\mathbf{\theta})$ on the entire dataset $\mathbf{x}$, we can compute $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x}_{b};\mathbf{\theta})$ on a minibatch of data $\mathbf{x}_{b}$ of size $b$ to perform stochastic optimization (see line 7 of Algorithm 1). The stochastic estimate provides an unbiased estimator of the gradient. In this way, SoftKI can leverage powerful stochastic optimization techniques used to train neural networks such as Adam (Kingma & Ba, 2014), and hardware acceleration such as graphics processing units (GPUs). The time complexity of evaluating $\nabla\log\hat{p}(\mathbf{y}\,|\,\mathbf{x}_{b};\mathbf{\theta})$ is cheap, costing $\mathcal{O}(b^{2}m)$ operations with space complexity $\mathcal{O}(bm)$. The lower space complexity makes it easier to use GPUs since they have more memory constraints compared to CPUs. We emphasize that we are performing stochastic gradient descent and not stochastic variational inference as in SVGP (Hensman et al., 2013). In particular, we do not define a distribution on the interpolation points $\mathbf{z}$ nor make a variational approximation.
@@ 3.2 Learning Interpolation Points
5: $\triangleright$ Model Training
[…]
@@
12:   $\mathbf{\theta}\leftarrow\mathbf{\theta}+\eta\,\nabla_{\mathbf{\theta}}\log\widehat{p}\Big{(}\mathbf{y}_{b}\,\big{|}\,\mathbf{x}_{b};\,\mathbf{K}_{\mathbf{x}_{b}\mathbf{x}_{b}}^{\text{SoftKI}}+\mathbf{\Lambda}\Big{)}$ $\triangleright$ Stabilized MLL (Section 3.2)
@@
15: $\triangleright$ Stabilized Inference (Section 3.3
[…]
@@ 3.3 Posterior Inference
$p(f(*)\,|\,\mathbf{x},\mathbf{y})=\mathcal{N}(\mathbf{\widehat{K}_{*\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{y},\mathbf{K}_{**}^{\text{SoftKI}}-\mathbf{K}_{*\mathbf{x}}^{\text{SoftKI}}(\mathbf{\Lambda}^{-1}-\mathbf{\Lambda}^{-1}\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}\mathbf{\widehat{C}}^{-1}\mathbf{\widehat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1})\mathbf{K}_{\mathbf{x}*}^{\text{SoftKI}})$
@@
for weights $\alpha$. Note that this is the posterior mean of a SGPR with $\mathbf{C}$ replaced with $\widehat{\mathbf{C}}$ and $\mathbf{K_{\mathbf{x}\mathbf{z}}}$ replaced with $\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}$. Moreover, note that $\mathbf{\widehat{C}}$ is simply $\mathbf{C}$ with $\mathbf{K_{\mathbf{x}\mathbf{z}}}$ replaced with $\mathbf{\widehat{K}_{\mathbf{x}\mathbf{z}}}$. Since $\mathbf{\widehat{C}}$ is a $m\times m$ matrix, the solution of the system of linear equations has time complexity $\mathcal{O}(m^{3})$. The formation of $\mathbf{\widehat{C}}$ requires the multiplication of a $m\times n$ matrix with a $n\times m$ matrix which has time complexity $\mathcal{O}(nm^{2})$. Thus the complexity of SoftKI posterior mean inference is $\mathcal{O}(nm^{2})$ since it is dominated by the formation of $\mathbf{\widehat{C}}$. We refer the reader to the supplementary material for discussion of the posterior covariance (Appendix A).
Solving with QR. Unfortunately, solving Equation 9 for $\alpha$ can be numerically unstable. Foster et al. (2009) introduce a stable QR solver approach for a Subset of Regressors (SoR) GP (Smola & Bartlett, 2000) which we adapt to a SoftKI. Since the form of the SoftKI and SGPR posterior are similar, this QR solver approach can also be adapted to improve the performance of SGPR (see Appendix B).
Define the block matrix
$\mathbf{A}=\begin{pmatrix}\mathbf{\Lambda}^{-1/2}\mathbf{\widehat{K}}_{\mathbf{x}\mathbf{z}}\\ \mathbf{U}_{\mathbf{z}\mathbf{z}}\end{pmatrix}$
@@
$\mathbf{A}^{\top}\mathbf{A}=\widehat{\mathbf{K}}_{\mathbf{z}\mathbf{x}}\mathbf{\Lambda}^{-1}\widehat{\mathbf{K}}_{\mathbf{x}\mathbf{z}}+\mathbf{K}_{\mathbf{z}\mathbf{z}}=\widehat{\mathbf{C}}$.
@@
We thus solve Equation 10 for $\alpha$ via a triangular solve in $\mathcal{O}(m^{2})$ time. For additional experiments demonstrating the inability of other linear solvers to offer comparable accuracy to the QR-stabilized solve detailed in this section, see Appendix B.